A rower of mass 70.0kg steps out of s stationary boat with a velocity of 2.50ms-1 forwards onto the nearby riverbank. the boat has a mass of 400kg and was initially at rest. with what velocity does the boat begin to move as the rower steps out? Give the answer to three significant figures.


1
Expert's answer
2021-08-10T08:00:56-0400

The initial momentum is zero, so, the final momentum must be conserved:


"p_i=p_f,\\
0=m_rv_r-m_bv_b,\\\space\\
v_b=\frac{m_rv_r}{m_b}=\frac{70·2.5}{400}=0.438\text{ m/s}."